Two cars are involved in a collision. If the accident sends the person in Car A heading toward the dashboard with an initial velocity of 30 m/s, and they will end with a final velocity of 0 m/s, and if the airbag deployed allows them to slow down over a space of 0.5 m, What will their acceleration be as they come to a stop?

A. 0.03 m/s^2
B. 30 m/s^2
C. 900 m/s^2
D. 0.001 m/s^2

A man on a high dive jumps up off of the platform with a velocity of + 3 m/s. If the platform is 4 meters above the water, with what velocity will the man hit the water?

A. - 9.35 m/s
B. - 87.48 m/s
C. - 11.3 m/s
D. -0.11 m/s

Part 2: What it the total time he spends in the air?

A. 0.65 seconds
B. 6.35 seconds
C. 0.31 seconds
D. 0.33 seconds
E. 1.26 seconds
F. 0.95 seconds

A car drives down the road at 50 km/hr. If it drives 80 km, how long (what amount of time) was it driving for?

A. 1 hr
B. 1.6 hrs
C. 1.5 hrs
D. 40 hrs
E. 0.625 hrs
1
Expert's answer
2020-10-13T09:29:22-0400

1)


"v^2_f-v_i^2=-2as\\0^2-30^2=-2(0.5)a\\a=900\frac{m}{s^2}"

Answer: C

2)


"v^2_f-v_i^2=2gh\\v_f^2-3^2=2(9.8)4\\v_f=-9.35\frac{m}{s}"

Answer: A

3)


"t=t_1+t_2=\frac{v_i}{g}+\frac{-v_f}{g}\\t=\frac{3}{9.8}+\frac{9.35}{9.8}=1.26\ s"

Answer: E

4)


"t=\frac{s}{v}=\frac{80}{50}=1.6\ hrs"

Answer: B